Of 11 possible; books, you plan to take 8 with you on vacation. How many different collections of 8 books can you; take?
Answer by stanbon(75887)

Of 11 possible; books, you plan to take 8 with you on vacation. How many different collections of 8 books can you; take?
-----
Note:: Choosing 8 of 11 is the same as not choosing 3 of 11.
Ans: 11C8 = 11C3 = (11*10*9)/(1*2*3) = 11*5*3 = 11*15 = 165
-------------
Cheers,
Stan H.
------------